Aparticle is projected upwards . The times according to height h while ascending and descending are 4s and 6s respectively . Find velocity of projection

Grade:11

1 Answers

ATUL TAMRAKAR
6 years ago
First let us assume initial velocity is u and height h then from equation of motion,t =4s,g=10m/sS= ut -1/2(gt^2). (For upward)h=u×4-1/2(10×(4)^2)u=(80+h)/2Then for descending,u=0,t=6sS= ut +1/2(gt^2). (For downward)h=0+1/2(10(6)^2)h=180Putting this in previous equation,we getu=(180+80)/4u=260/4u=65Hence this is the velocity of projection.